\begin{thebibliography}{aaaaa}

\bibitem{A2}
L. Ambrosio. \newblock Some fine properties of sets of finite perimeter in Ahlfors regular metric measure spaces.
\newblock {\em Advances in Mathematics} 159 (2001), 51--67.

\bibitem{AFP}
L. Ambrosio, N. Fusco, and D. Pallara.
\newblock {\em Functions of bounded variation and free discontinuity problems}.
\newblock Oxford Mathematical Monographs. The Clarendon Press, Oxford
  University Press, New York, 2000.

\bibitem{ADS}
L. Ambrosio, S. Di Marino and G. Savar\'e. \newblock 
On the duality between p-modulus
and probability measures.
\newblock {\em J. Eur. Math. Soc.} 17 (2015), 1817-1853.

\bibitem{AMP}
L. Ambrosio, M. Miranda Jr., and D. Pallara. \newblock{
Special functions of bounded
variation in doubling measure spaces.} 
\newblock{ \em Calculus of variations: topics
from the mathematical heritage of E. De Giorgi,} 1--45, 2004.


\bibitem{BB}
A. Bj{\"o}rn and J. Bj{\"o}rn. \newblock {\em Nonlinear Potential theory on metric spaces}. \newblock  EMS Tracts in Mathematics,
17. European Mathematical Society (EMS), Zurich, 2011.

\bibitem{Br} M. Brelot.
\newblock{\em Lectures on potential theory.} 
\newblock{Notes by K. N. Gowrisankaran and M. K. Venkatesha Murthy. Second edition, revised and enlarged with the help of S. Ramaswamy. 
Tata Institute of Fundamental Research Lectures on Mathematics, No. 19 Tata Institute of Fundamental Research, Bombay, 1967}


\bibitem{Cho} G. Choquet. \newblock Theory of capacities.
\newblock {\em Annales de l'Institut Fourier} 5 (1953--54), 131--295.


\bibitem{EG}
L.~C. Evans and R.~F. Gariepy.
\newblock {\em Measure theory and fine properties of functions}.
\newblock Studies in Advanced Mathematics. CRC Press, Boca Raton, FL, 1992.

\bibitem{Fed}
H.~Federer.
\newblock {\em Geometric measure theory}.
\newblock Die Grundlehren der mathematischen Wissenschaften, Band 153.
  Springer-Verlag, New York, 1969.
  
\bibitem{Fle}
W. H. Fleming.
\newblock
Functions whose partial derivatives are measures.
\newblock{\em Illinois J. Math.} 4 (1960), 452--478.

\bibitem{Fug}
B.~Fuglede.
\newblock Extremal length and functional completion.
\newblock {\em Acta Math.} 98 (1957), 171--219.

\bibitem{HKM}
 J. Heinonen, T. Kilpel\"ainen and O. Martio. \newblock {\em Nonlinear potential
 theory of degenerate elliptic equations.} \newblock  Dover Publications Inc., New York, 2006.
 



\bibitem{HKST}
 J. Heinonen, P. Koskela, N. Shanmugalingam
and J. Tyson. \newblock {\em Sobolev spaces on metric measure spaces.} \newblock  Cambridge University 
Press, 2015.





\bibitem{HEKMM}
V.~Honzlov\'a~Exnerov\'a, O.~Kalenda, J.~Mal\'y and O.~Martio.
\newblock Plans on measures and $AM$--modulus.
\newblock {To appear}

\bibitem{HEMM}
V.~Honzlov\'a~Exnerov\'a, J.~Mal\'y and O.~Martio.
\newblock Modulus in {B}anach function spaces.
\newblock {\em Ark. Mat.} 55(1) (2017), 105--130.

\bibitem{HEMM2}
V.~Honzlov\'a~Exnerov\'a, J.~Mal\'y and O.~Martio.
\newblock {Functions of bounded variation and the ${AM}$--modulus in $\rn$.}
\newblock {\em Nonlinear Anal.} 177 (2018), 553--571. 

\bibitem{Kech} 
A. S. Kechris.
\newblock {\em Classical descriptive set theory.}
\newblock {Graduate Texts in Mathematics, 156. Springer-Verlag, New York, 1995.}

\bibitem{KKST1}
J. Kinnunen, R. Korte, N. Shanmugalingam and H. Tuominen.
\newblock {Lebesgue points and capacities via the boxing inequality in metric spaces.}
\newblock {\em Indiana Math. J.}
 57, no. 1 (2008), 41--67.

\bibitem{KKST}
J. Kinnunen, R. Korte, N. Shanmugalingam and H. Tuominen.
\newblock {Pointwise properties of functions of bounded variation in metric spaces.}
\newblock {\em Rev. Mat. Complut.} 27 (2014), 41--67.



\bibitem{KL}
R. Korte and P. Lahti.
\newblock {Relative isoperimetric inequalities and sufficient conditions for finite perimeter on metric spaces.}
\newblock {\em Annales de l'Institut Henri Poincar\'e Non Linear Analysis}
 31(1) (2014), 129--154.

\bibitem{La}
P. Lahti. 
\newblock {A new Federer-type characterization of sets of finite perimeter in metric spaces.}  
\newblock {Preprint, to appear.}

\bibitem{M}
O. Martio. \newblock Functions of bounded variation and curves in metric measure spaces. \newblock {\em  Adv. Calc. Var.}  Vol. 9(4) (2016), 305-322.  


\bibitem{M1}
O. Martio. \newblock  
The space of functions of bounded variation on curves in metric measure spaces. \newblock {\em
Conform. Geom. Dyn.} 20 (2016), 81--96.

\bibitem{Mi}
M. Miranda Jr. \newblock Functions of bounded variation on "good" metric spaces. \newblock {\em J. Math. Pures Appl.} (9) 82 (2003), 975--1004.

\bibitem{Roy} 
H. L. Royden. \newblock {\em Real analysis.} \newblock  Third edition. Macmillan Publishing Company, New York, 1988.

\end{thebibliography}

\documentclass[a4paper,draft]{amsart}
\usepackage{amsthm,enumerate,
verbatim,amssymb,color}

\def\Cap{\operatorname{Cap}}
\def\Cp{C_*}

\def\E{\mathcal E}

\def\H{\mathit{co}\mathcal H}

\def\K{\mathcal K}

\def\diam{\operatorname{diam}}
\def\dist{\operatorname{dist}}

\def\er{\mathbb R}

\def\rn{\mathbb R^n}

\def\haus{\mathcal H^{n{-}1}}

\def\er{\mathbb R}
\def\Lip{\operatorname{Lip}}
\def\cp{\operatorname{cap}}
\def\Cp{\operatorname{Cap}}
\def\loc{\mathrm{loc}}

\def\rn{\mathbb R^n}

\def\inf{\operatorname{inf}}
\def\sup{\operatorname{sup}}

\newdimen\vintkern
\newdimen\vintkern
\vintkern10pt\usepackage{amsmath}
\usepackage{amsfonts}
\usepackage{amssymb}

\def\vint{{-}\kern-\vintkern\int}

\def\eqn#1$$#2$${\begin{equation}\label#1#2\end{equation}}
\def\inte{\operatorname{int}_*\!}
\def\exte{\operatorname{ext}_*\!}

\def\standhaus{\widetilde{\mathcal H}^{n{-}1}}

\def\H{\mathcal H}

\newtheorem{thm}{Theorem}[section]
\newtheorem{lemma}[thm]{Lemma}
\newtheorem{cor}[thm]{Corollary}
\newtheorem{prop}[thm]{Proposition}

\theoremstyle{definition}

\newtheorem{examples}[thm]{Examples}

\newtheorem{rmrk}[thm]{Remark}

\long\def\colred#1\endred{{\color{red}#1}}
\long\def\colgreen#1\endgreen{{\color{green}#1}}
\long\def\colblue#1\endblue{{\color{blue}#1}}

\title[$AM$--modulus and and Hausdorff measure of codimension one]{\vspace{0.5in}
{\bf $AM$--modulus and Hausdorff measure of codimension one in metric measure spaces\\}}

\author{Vendula Honzlov\'a Exnerov\'a}
\address{Department of Mathematical Analysis, Faculty of Mathematics and Physics,
Charles University in Prague, So\-ko\-lovsk\'a 83,  Prague 8, 186\,75 Czech Republic}
\email{honzlova@karlin.mff.cuni.cz}
\author{Jan Mal\'y}
\address{Department of Mathematical Analysis, Faculty of Mathematics and Physics,
Charles University in Prague, So\-ko\-lovsk\'a 83,  Prague 8, 186\,75 Czech Republic}
\email{maly@karlin.mff.cuni.cz}

\author{Olli Martio}
\address{Department of Mathematics and Statistics, FI-00014 University of Helsinki, Finland}
\email{olli.martio@helsinki.fi}

\thanks{The first and the second authors have been supported by the grant GA\,\v{C}R 
P201/18-07996S of the Czech Science Foundation.}

\begin{document}

\begin{abstract}
Let $\Gamma(E)$ be the family of all paths which
meet a set $E$ in the metric measure space $X$. The set function $E \mapsto AM(\Gamma(E))$ defines the $AM$--modulus measure in $X$ where $AM$ refers to the approximation modulus \cite{M}. We 
  compare $AM(\Gamma(E))$ to the Hausdorff measure $co\H^1(E)$ of codimension one in $X$ and show that
$$co\H^1(E) \approx AM(\Gamma(E))$$
 for Suslin sets $E$ in $X$. This leads to  a new characterization of sets of finite perimeter in $X$ in terms of the $AM$--modulus. We
 also study the level sets of $BV$ functions and show that for a.e. $t$ these sets
 have finite $co\H^1$--measure.
Most of the results are new also in $\rn$.
\end{abstract}

\keywords
{$AM$--modulus; perimeter; sets of co--dimension one; level sets of $BV$-functions; metric measure spaces}

\maketitle

\section{Introduction}

In a metric measure space $X$ the modulus of a curve
family offers a substitute for the Fubini theorem
and provides an important tool for analysis in $X$, see e.g. \cite{BB}. The $M_p$--modulus,
$p \geq 1$, is used to create a space  in $X$ similar to the
Sobolev space $W^{1,p}(\rn)$ and the $AM$--modulus
 was introduced as a weaker version than the
$M_1$--modulus to study functions of bounded variation in $X$ and in $\rn$, see \cite{M}, \cite{HEMM} and \cite{HEMM2}.

Let $\Gamma(E)$ be the family of all
paths in $X$ which meet the set $E \subset X$.
The set function $E \mapsto AM(\Gamma(E))$ defines a metric outer measure, the $AM$--modulus measure, in $X$ and satisfies
\begin{equation} \label{Eq1}
AM(\Gamma(E)) \leq C\, co\H^1(E)
\end{equation}
provided that the measure $\mu$ is doubling, see
Theorem \ref{Tupper} below. Here $co\H^1$ refers to
the Hausdorff measure of codimension one in $X$. We also present the generalization of
(\ref{Eq1}) for all measures $co\H^p$, $p \geq 1$.

In this paper we are interested in the inequalities opposite to (\ref{Eq1}). Such an 
inequality was obtained in \cite{HEMM2} for sets $E$ contained in $(n-1)$--rectifiable 
sets in $\rn$. Here we show that this inequality holds for 
Borel sets in $X$, and more generally for Suslin sets and for arbitrary 
sets with $\sigma$--finite $co\H^1$--measure, provided that $X$ satisfies standard regularity assumptions, i.e. the measure $ \mu$ in $X$ is doubling, $X$ is complete and supports the
Poincar\'e inequality.
 Thus in $\rn$ the standard $(n-1)$--Hausdorff measure $\H^{n-1}$ satisfies
\begin{equation} \label{EHRn}
 \H^{n-1}(E) \approx AM(\Gamma(E))
 \end{equation}
for all Suslin sets and arbitrary sets of $\sigma$--finite $\H^{n-1}$--measure. Note that the
ordinary $M_p$--modulus is more adapted to measure
the family $\Gamma(E, \Omega)$ of all curves which join $E$ to the complement of a fixed open set $\Omega$ and then the result corresponds to the
$p$--capacity of $E$. Thus the relation to
the $(n-p)$--dimensional Hausdorff measure is mediated through the capacity and does not provide as close 
a connection as (\ref{EHRn}), see also Remark \ref{RAMRn}.

We apply the above results to study the $AM$--modulus of path families which are closely associated with sets of finite perimeter in $X$. 
Although there is extensive literature on sets of finite perimeter in metric measure spaces, see \cite{A2}, \cite{AFP}, \cite{KKST},
\cite{KL}, \cite{La} and \cite{Mi}, the $AM$--modulus has not yet been used to characterize sets of finite perimeter in $X$ and our results extend the characterizations obtained in \cite{HEMM2} 
in $\rn$ to $X$. 

We study  the level sets of a $BV$ function $u$ in the final section and show that these sets
have finite $co\H^1$--measure for a.e. $t$. In particular, it follows that the ordinary level 
set $u^{-1}(t)$ of a continuous $BV$ function 
$u$  has finite $co\H^1$--measure for a.e. $t$.

\section{Preliminaries}

Let $(X,d)$ be a metric space and $\mu$ a Borel regular measure in $X$. The measure 
$\mu$ is \textit{doubling} if there is a constant $C_{\mu}$ such that 
$\mu(B(x,2r)) \leq C_{\mu}\, \mu(B(x,r))$ and $0 <\mu(B(x,r))<\infty$ for 
all open balls $B(x,r)$ in $X$.

A continuous mapping $\gamma\colon [a,b]\to X$ is called a \textit{curve}.
We say that a curve $\gamma$ is a \textit{path}
if it has a finite and non--zero total length; in this case we parametrize
$\gamma$ by its arclength. 
The \textit{locus} of $\gamma$ is defined as $\gamma([0,\ell])$ and 
denoted by $\langle\gamma\rangle$.

We refer to \cite{M} and \cite{HEMM} for the properties of the $AM_p$--modulus and to 
\cite{BB} and \cite{Fug} for those of the $M_p$--modulus. For completeness we recall the definitions.

Let $\Gamma$ be a family of paths in $X$. 
A non--negative Borel function $\rho$ is $M$--\textit{admissible}, or simply admissible, for $\Gamma$ if
$$ \int_{\gamma} \rho \, ds \geq 1$$
for every $\gamma \in \Gamma$. For $p \geq 1$ the $M_p$--\textit{modulus} of $\Gamma$ is defined as
$$M_p(\Gamma) = \inf \int_X \rho^p \, d\mu$$
where the infimum is taken over all admissible functions $\rho$. 

A sequence of non--negative Borel functions $\rho_i $, $i = 1, \,2, \,...\,,$ is 
$AM$--\textit{admissible}, or simply admissible,
for $\Gamma$ if
\begin{equation}
\label{eqn1}
\liminf_{i \rightarrow \infty} \int_{\gamma} 
\rho_i \, ds \geq 1
\end{equation}
for every $\gamma \in \Gamma$. The \textit{approximation
modulus}, $AM_p$--modulus for short, of $\Gamma$ is defined as
\begin{equation}
\label{eqn2}
AM_p(\Gamma) = \inf_{(\rho_i)} \Big\{ \liminf_{i \rightarrow \infty} \int_{X} \rho_i^p \, d\mu \Big\}
\end{equation}
where the infimum is taken over all $AM$--admissible sequences $(\rho_i)$ for $\Gamma$. We mostly consider the $AM_1$--modulus and use the abbreviation $AM = AM_1$. Note that for $p> 1$, $AM_p(\Gamma) = M_p(\Gamma)$ for every path family $\Gamma$ in $X$, see \cite[Theorem 1]{HEMM}, however, sometimes it is easier to use the $AM_p$--modulus than the $M_p$--modulus.
Note also that $AM(\Gamma)\le M_1(\Gamma)$ for all path families $\Gamma$ in $X$ and
it could happen that $AM(\Gamma)= 0$ but
$M_1(\Gamma) = \infty$ for some family $\Gamma$.

We define the $AM_c$--modulus of $\Gamma$ with respect to the $AM$--modulus with the 
difference that the admissible sequence are now required to consist of 
continuous functions.

The $AM$ modulus or the $AM_c$ modulus can be also assigned to a 
family $\mathcal E$ of measures, $\int_{\gamma}\rho_i\,ds$, 
$\gamma\in\Gamma$, is then 
replaced by $\int_X\rho_i\,d\nu$, $\nu\in \mathcal E$. For a more 
precise definition we refer to \cite{HEKMM}. 

For $E \subset X$,  $\Gamma(E)$ denotes the family of all paths which meet $E$. From \cite[Theorem 1]{HEMM2} it follows that 
the set function $\phi: E \mapsto AM(\Gamma(E))$ 
is a metric outer measure in $X$ and hence all Borel sets are $\phi$ measurable. Almost the same proof shows that for $p \geq 1$ the set functions $E \mapsto AM_p(\Gamma(E))$
and  $E \mapsto M_p(\Gamma(E))$ also define metric outer measures in $X$. 

We denote by $\H^{n-p}$ the ordinary Hausdorff measure
of codimension $p$ in $\rn$. In metric spaces, the dimension $n$ is not always clearly determined. The right replacement of $\H^{n-p}$ is then the \textit{Hausdorff
measure $co\H^p(E)$ of codimension} $p$ defined as
$$co\H^p(E)=\sup_{\delta>0}\, co\H^p_{\delta}(E)$$
where
for $\delta > 0$
$$co\H^p_{\delta}(E)=\inf
\Bigl\{
\sum_{j=1}^{\infty}\frac{\mu(B(x_j,r_j))}{r_j^p}\colon E\subset
\bigcup_{j=1}^{\infty} B(x_j,r_j),\;\sup_jr_j<\delta
\Bigr\}$$
denotes the $\delta$--content associated with $ co\H^p(E)$. It is easily checked that in $\rn$, $co\H^p$ agrees with the $\H^{n-p}$--measure up to a multiplicative constant. 

In the following, we are chiefly interested in $co\H^1(E)$ and its dependence on $AM(\Gamma(E))$ and we first consider upper bounds for $AM(\Gamma(E))$ in terms of $co\H^1(E)$. Such a
result was presented in
\cite[Theorem 3.17]{M} and for completeness we include a proof. For $p>1$ we present a stronger version in $X$ and extend the implication,  see \cite[Theorem 2.27]{HKM} and references therein, that in $\rn$,
$\H^{n-p}(E) < \infty$ implies that the $p$--capacity of $E \subset \rn$ is zero.

\begin{thm} \label{Tupper}
Suppose that $\mu$ is a doubling measure in $X$ and $E \subset X$. Then 
\begin{equation}  \label{ETu1}
AM(\Gamma(E)) \leq C_{\mu} \, co\H^1(E)
\end{equation}
and for $p > 1$, 
$co\H^p(E)<\infty$ implies 
$M_p(\Gamma(E))=0$. 
\end{thm}

\begin{proof} 
First, we prove
\begin{equation}  \label{ETup} 
AM_p(\Gamma(E)) \leq C_{\mu} \, co\H^p(E)
\end{equation}
for any $1\le p<\infty$.
We may assume that $co\H^p(E) < \infty$. 
For $j = 1, \,2,\, ...$ choose a covering $B(x^j_i,r^j_i), \,
i=1,\,2,\, ...$ , of $E$ such that $r^j_i < 1/j$ and
$$ \sum_i \frac{\mu(B(x^j_i,r^j_i))}{(r_i^j)^p} \leq co\H^p_{1/j}(E)  +\frac{1}{j}.$$
Set
$$\rho_j(x) = \Bigl\{ \sum_i\frac
{1}{(r^j_i)^p}\chi_{B^j_i}(x)\Bigr\}^{1/p}$$
where $B^j_i = B(x^j_i,2r^j_i)$.
Then $\rho_j$ is a Borel function and we show that the sequence $(\rho_j)$ is admissible for 
$\Gamma(E)$. Indeed, if $\gamma \in \Gamma(E)$, then
$\gamma$ meets $E$ and since $\gamma$ is not a constant path, $\diam \, \langle\gamma\rangle > 4/j$ for large $j$
and hence there is $j_0$ such that for $j \geq j_0$ we find $i = i(j)$ such that $\gamma$ meets $B(x^j_i, r^j_i)$
and $X \setminus B^j_i$. Thus $\gamma$ travels in $B^j_i$ at least distance $r^j_i$.
Consequently for $j \geq j_0$
$$ \int_{\gamma} \rho_j \, ds \geq \int_{\gamma } \frac{ \chi_{B^j_{i(j)}} }{r^j_{i(j)}}   \, ds \geq 1$$
and hence
$$\liminf_{j \rightarrow \infty}\int_{\gamma} \rho_j \, ds \geq 1.$$
We obtain
$$ 
\aligned
AM_p(\Gamma(E)) &\leq \liminf_{j \rightarrow \infty}\int_{X} \rho_j^p \, d\mu 
= \liminf_{j \rightarrow \infty}\, \sum_i\frac
{\mu(B^j_i)}{(r^j_i)^p} 
\\&
 \leq C_{\mu} \liminf_{j \rightarrow \infty}\sum_i \frac{\mu(B(x^j_i,r^j_i)) }{(r^j_i)^p} 
\leq C_{\mu} \liminf_{j \rightarrow \infty}\, ( co\H^p_{1/j}(E) + \frac{1}{j}) 
\\&=C_{\mu} \, co\H^p(E),
\endaligned
$$
which proves \eqref{ETup}

Now, for $p=1$ we are done. 
If $p>1$, we know by \cite[Theorem 1]{HEMM}
that $M_p=AM_p$, therefore we have
\eqn{ETupa}
$$
M_p(\Gamma(E)) \leq C_{\mu} \, co\H^p(E).
$$
To prove that $M_p(\Gamma(E))=0$, we first use 
\eqref{ETupa} to
construct a sequence $(\rho_j)$ of $M$--admissible functions for $\Gamma(E)$
such that 
\eqn{CCp}
$$
\int_{X}\rho_j^p\,d\mu\le C\text{ with }C=1+C_{\mu} \, co\H^p(E)
$$ 
and $\mu(\{\rho_j>0\})\to 0$. 
Note that  $\mu(\{\rho_j>0\})$ can be made arbitrary small. To see this let $\varepsilon > 0$ and since $\mu(E) = 0$ we can choose an open
set $G \supset E$ with $\mu(G) < \varepsilon$. 
If $\rho$ is admissible for $\Gamma(E)$, we set 
$$
\tilde\rho=
\begin{cases}
\rho&\text{in }G,\\
0&\text{ in }X\setminus G.
\end{cases}
$$
Each path $\gamma \in \Gamma(E)$
has a subpath $\tilde{\gamma}\in \Gamma(E)$ with locus in $G$. Then
$$
\int_{\gamma}\tilde\rho\,ds \ge \int_{\tilde\gamma}\rho\,ds\ge 1,
$$
and thus $\tilde \rho$ is admissible for $\Gamma(E)$ as well.
Moreover, $\mu(\{ \tilde{\rho}> 0 \}) < \varepsilon$ and
$$ \int_X \tilde{\rho}^p \, d\mu \leq \int_X \rho^p \, d\mu .$$

Now, we select a special subsequence.
We proceed by induction.  Set $m_1=1$. If $m_1,\dots,m_{j-1}$ are determined,
we find $m_j$ such that 
\eqn{2-j}
$$
\int_{E_j}(\rho_{m_1}+\dots+\rho_{m_{j-1}})^p\,d\mu<2^{-j},
$$
holds with $E_j=\{\rho_{m_j}>0\}$.
We claim that 
\eqn{sump}
$$
\int_X(\rho_{m_1}+\dots+\rho_{m_{j}})^p\,d\mu\le 2^{p-1}(Cj+1).
$$
Indeed, it follows from \eqref{CCp} as we prove
\eqn{induc}
$$
\int_X(\rho_{m_1}+\dots+\rho_{m_{j}})^p\,d\mu\le
2^{p-1}\Bigl(\int_X(\rho_{m_1}^p+\dots+\rho_{m_{j}}^p)\,d\mu+\sum_{i=1}^{j}2^{-i}\Bigr)
$$
by induction.
The inequality is trivial for $j=1$.
If it holds for $j-1$, using  \eqref{2-j} we obtain
$$
\aligned
\int_{X}(\rho_{m_1}&+\dots+\rho_{m_{j}})^p\,d\mu
\le \int_{X\setminus E_j}(\rho_{m_1}+\dots+\rho_{m_{j-1}})^p\,d\mu
\\&\qquad+\int_{E_j}(\rho_{m_1}+\dots+\rho_{m_{j}})^p\,d\mu
\\&
\le 2^{p-1}\Bigl(\int_X(\rho_{m_1}^p+\dots+\rho_{m_{j-1}}^p)\,d\mu+\sum_{i=1}^{j-1}2^{-i}\Bigr)
\\&\qquad+2^{p-1}\Bigl(\int_X\rho_{m_j}^p\,d\mu+\int_{E_j}(\rho_{m_1}+\dots+\rho_{m_{j-1}})^p\,d\mu\Bigr)
\\&
\le 2^{p-1}\Bigl(\int_X(\rho_{m_1}^p+\dots+\rho_{m_{j}}^p)\,d\mu+\sum_{i=1}^{j}2^{-i}\Bigr)
\endaligned
$$
which proves \eqref{induc} for $j$.

Finally, we test the $M_p$-modulus of $\Gamma(E)$ by the admissible functions
$$
g_k=\frac{1}k\sum_{j=1}^k\rho_{m_j}.
$$
Then it is evident that each $g_k$ is admissible for $\Gamma(E)$ and 
by \eqref{sump} 
$$
M_p(\Gamma(E)) \le \int_{X}g_k^p\,d\mu\le 2^{p-1}k^{-p}(Ck+1).
$$
\end{proof}

\begin{rmrk}
\label{RAMRn}
Consider the inverse implication in Theorem \ref{Tupper} for $p > 1$ in $\rn$.
Let $E \subset \rn$ be a Borel set with
$M_p(\Gamma(E)) < \infty$, $1 < p \le n$.
If $K \subset E$ is compact, then
$$M_p(\Gamma(K)) \le M_p(\Gamma(E)) < \infty$$
and it easily follows that for all open sets
$\Omega \supset K$
$$ \cp_p(K, \Omega) \le  M_p(\Gamma(K))$$
where $ \cp_p(K, \Omega)$ stands for the
ordinary variational $p$--capacity of the
condenser $(K, \Omega)$, see Section 3 and \cite[Chapter 2]{HKM}.
From \cite[Lemma 2.34]{HKM} it follows that $K$ has $p$--capacity zero and hence by the Choquet capacitability theorem $E$ has also capacity zero. This implies, see e.g. \cite[Theorem 2.27]{HKM}, that
the Hausdorff dimension of $E$ is at most $n-p$ but not that $\H^{n-p}(E) < \infty$.
\end{rmrk}
 
We also need some properties of functions of bounded
variation ($BV$) in $X$, see \cite{Mi} (in metric measure spaces) and 
\cite{AFP} (in the Euclidean spaces). Let $\Omega \subset X$ be open and
denote by $\Lip_{\textnormal{loc}}(\Omega)$ the set of locally Lipschitz
functions in $\Omega$.
Given $u \in L^1_{\textnormal{loc}}(\Omega)$ and an open set $G\subset \Omega$ we 
define
$$ 
V(u,G) = \inf
\Bigl\{ \liminf_i \int_{G} |\nabla u_i| \, d\mu \colon u_i \rightarrow u \textnormal{ in 
} L^1_{\loc}(G)\Bigr\}
$$
Here $|\nabla u(x)|$ stands for the local Lipschitz constant for $u$ at $x$, i.e.
$$ |\nabla u(x)| =  \liminf_{r \rightarrow 0} \sup_{y \in B(x,r)} \frac{|u(y) - u(x)|}{r},$$
see \cite[Section 1.3]{BB}.
A function  has 
\textit{bounded variation} in $\Omega$, $u \in BV(\Omega)$, if 
$V(u,\Omega)<\infty$.

Let $\Omega \subset X$ be open and let $E \subset X$ be measurable. 
The \textit{perimeter} of $E$ in $\Omega$ is 
$P(E,\Omega)=V(\chi_E,\Omega)$ and we write
$P(E)=P(E,X).$

The space $X$ supports the (weak) $BV$--Poincar\'e inequality, see \cite[Remark 3.5]{Mi}, if
\begin{equation} \label{EPBV}
 \int_{B(x,r)}|u -u_{B(x,r)} |\, d\mu \leq C_P \, r\,V(u,B(x,\lambda_P r))
\end{equation}
in each ball $B(x,r)$ and for each $u \in BV(X)$. Here $u_{B(x,r)}$ stands for 
the mean value of $u$ in $B(x,r)$. The constants $C_P\geq 1$ and 
$\lambda_P \geq 1$ are independent of $B(x,r)$ and $u$ and called the Poincar\'e
constants of $X$. Note that (\ref{EPBV}) is a consequence of 
the standard weak Poincar\'e inequality for integrable 
functions with upper gradients, see \cite[Chapter 4]{BB} and \cite{Mi}. 

We use the standard assumptions (A) on the space $X$:
\begin{itemize}
\item $X$ is complete,
\item the measure $\mu$ is doubling,
\item $X$ supports the $BV$--Poincar\'e inequality
(\ref{EPBV}).
\end{itemize}
Note that if $\mu$ is doubling and $X$ is complete,
then $X$ is proper, i.e. closed and bounded subsets of
$X$ are compact, see \cite[Section 3.1]{BB}. Moreover, $X$
is connected \cite[Proposition 4.2]{BB}.

\section{Newtonian and perimeter capacities in $X$}

Throughout this and the next section we assume that $(X,d)$ and $\mu$ satisfy the assumptions (A). 

Let $G$ be a bounded open set in $X$, let $K$ be a compact subset of $G$ and 
let $ \Lip_0(K,G) $ be the set of all Lipschitz functions $u$ with compact support in $G$
satisfying  $u \geq 1$ on $K$. We define 
\eqn{capcomp}
$$ \cp_1(K,G) = \inf \Big\{
\int_G |\nabla u| \, d\mu: \, u \in \Lip_0(K,G) \Big\} .$$
Obviously the infimum does not change if restricted to test functions satisfying $0 \leq u \leq 1$.

It is easy to see that $\Lip_0(K,G) \neq \emptyset$
if $G  \neq \emptyset$ and thus $ \cp_1(K,G) < \infty$. Note that if $G$ is 
compact, then the constant function $1$ is a competitor and thus $ \cp_1(K,G)= 0$.

If $U \subset G$ is open, then we set
$$ \cp_1(U,G) = \sup \{ \cp_1(K,G): \,
K \subset U \textnormal{ compact}\}$$
and for an arbitrary set $E \subset G$
$$  \cp_1(E,G) = \inf \{ \cp_1(U,G): \,
 U \textnormal{ open }, E \subset U \subset G \}.$$ 

Now there are two definitions for $\cp_1(E,G)$ when $E$ is compact 
but since the competitors are continuous the next lemma is immediate. 

\begin{lemma}  \label{Lcoincide}
If $K \subset G$ is compact, then
\begin{equation} \label{Ecoin}
\cp_1(K,G) = \inf \{\cp_1(U,G) : \, U \textnormal{ open, }  K \subset U \subset G\},
\end{equation}
where the capacity on the left is according to \eqref{capcomp}.
\end{lemma}

Next we summarize the main properties of the capacity.
In particular, we show that $\cp_1(\cdot,G)$
defines a Choquet capacity and thus, by the  Choquet capacitability theorem, 
each Suslin (in particular, a Borel) set $E \subset G$ is capacitable.

We also compare  the widely used Newtonian type $p$--capacity 
\begin{equation} \label{ENew}
\widetilde{\cp}_p(E,G) =  \inf_u \int_G (g_u)^p \, d\mu
\end{equation}
for $p = 1$ to $\cp_1(E,G)$. In (\ref{ENew})
the infimum is taken over all (precisely defined)
$u \in N^{1,p}_0(G)$ such that $u \geq 1$ on $E$ and
$g_u$ is the minimal upper gradient of $u$, see
\cite[Section 6.3]{BB}. This is a Choquet capacity if
$p > 1 $ but not in the case $p=1$ because $\widetilde{\cp}_1$
does not satisfy (\ref{i:upwards}) below. For an example see 
\cite[Example 6.18]{BB} where it also becomes evident how
$\cp_1(E,G)$ differs from $\widetilde{\cp}_1(E,G)$.

\begin{prop}\label{p:cap}
\begin{enumerate}[\rm(a)]
\item\label{i:monotone} The set function $E \mapsto \cp_1(E,G)$ is monotone, i.e.
$$
E_1 \subset E_2 \subset G,
  \implies \cp_1(E_1,G) \le \cp_1(E_2,G).  
$$
\item\label{i:compact}
If $K_1,K_2,\dots\subset G$ are compact
and $K_1\supset K_2\supset\dots$,
then
$$
\cp_1\Big( \bigcap_{j=1}^{\infty}K_j , G\Big)
=\lim_{j\to\infty} \cp_1(K_j, G).
$$
\item\label{i:tilde}
$
\cp_1(E,G)\le \widetilde{\cp}_1(E,G)
\text{ and }\cp_1(K,G)= \widetilde{\cp}_1(K,G)
\text{ if }K\text{ is compact} .
$
\item\label{i:strong} 
If $K_1,K_2$ are compact, then 
$$
\cp_1(K_1\cup K_1,G)+\cp_1(K_1\cap K_2,G)\le \cp_1(K_1,G)+\cp_1(K_2,G).
$$
\item \label{i:upwards}
$
 E_1 \subset E _2 \subset \dots \subset G 
\implies \cp_1 \Big(\bigcup_{j=1}^{\infty}E_j,G \Big)
=\lim_{j\to\infty}\cp_1(E_j,G).
$
\item\label{i:Choquet} 
If $E\subset G$ is Suslin, then
$$
\cp_1(E,G) = 
\sup 
\{ \, \cp_1(K,G) : \, K \subset E \textnormal{ compact } \}.
$$
\end{enumerate}
\end{prop}

\begin{proof}
The properties \eqref{i:monotone} and \eqref{i:compact} are obvious.
The inequality in \eqref{i:tilde} is obvious if $E$ is open; for the
case of $E$ arbitrary we use \cite[Theorem 6.19 (vii)]{BB}
(note that the symbol $\cp_1$ stands for $\widetilde{\cp_1}$ in
\cite{BB}).
The equality for $K$ compact follows from \cite[Theorem 6.19 (x)]{BB}.
The property \eqref{i:strong} follows from \cite[Theorem 6.17 (iii)]{BB}
taking into account the equality in \eqref{i:tilde}.
Now, the properties \eqref{i:upwards} and \eqref{i:Choquet} are obtained
using the general theory of capacities developed by Choquet in \cite{Cho},
see also \cite{Br}, \cite{Kech}.
\end{proof}

If $G$ is a bounded open set in $X$ and $K \subset G$ compact, then we denote by $\Gamma(K,G)$ the family of all paths in $X$ which connect $X \setminus G$ to $K$.

\begin{lemma}\label{Lcap}
If $G$ is a bounded open set in $X$ and $K \subset G$ compact, then
$$
 \cp_1(K,G) = M_1(\Gamma(K,G)) =  AM(\Gamma(K,G)).
$$
\end{lemma}

\begin{proof}
Since for each function $u \in \textnormal{Lip}_0(K,G)$, $|\nabla u|$ is $M$--admissible
for the family $\Gamma(K,G)$, we
have $M_1(\Gamma(K,G)) \leq \cp_1(K,G)$. For the converse inequality we use the method in
\cite[Section 5.2]{BB}. Let $\rho$ be $M$--admissible for $\Gamma(K,G)$ and $\varepsilon > 0$. We may assume
that $\rho$ is lower semi--continuous.
From Lemmata 5.25 and 5.26 in \cite{BB} it follows that the function $\rho + \varepsilon$
is an upper gradient of the lower semi--continuous function 
$$ u(x) = \min \Big(1, \inf_{\gamma}  \int_{\gamma} (\rho +\varepsilon) \, ds \Big)$$
 in $G$. Here the infimum is taken over all paths connecting $X 
 \setminus G$ to $x \in G$. Moreover, $u = 0$ in $X\setminus G$ and $u = 1$ in $K$. 
Using  Proposition \ref{p:cap}\eqref{i:tilde} we obtain 
$$\cp_1(K,G) =\widetilde{\cp_1}(K,G)\leq \int_G (\rho +\varepsilon) \, d\mu \leq \int_G \rho \, d\mu 
+ \varepsilon\, \mu(G)$$
and letting $\varepsilon \rightarrow 0$ we obtain the desired inequality.

For the second equality it suffices to show that $M_1(\Gamma(K,G)) \leq  AM(\Gamma(K,G))$ because $M_1(\Gamma) \geq AM(\Gamma)$ for every path family $\Gamma$ in $X$. Let 
$\Gamma(K,G,L)$ denote the family of all the paths $\gamma$
in $\Gamma(K,G)$ whose length $\ell$ satisfies $ \ell \le L$. Note that 
\begin{equation}
\label{EM1}
M_1(\Gamma(K,G))=\sup_{L}M_1(\Gamma(K,G,L)).
\end{equation}
Indeed, if $\rho$ is admissible for $\Gamma(K,G,L)$, then
$\rho + \frac{1}{L}\chi_G$ is admissible for $\Gamma(K,G)$.

Fix $L$. Each $\gamma\in\Gamma(K,G,L)$ has a reparametrization $\xi\colon [0,L]\to X$
which is a curve with  $\Lip\xi\le1$; we denote the set of all such reparametrizations by $\Xi(K,G,L)$.
For a Borel set $E\subset X$ set
$$
\nu_{\xi}(E)=\int_{\xi}\chi_E\,ds.
$$
Set $\E=\{\nu_{\xi}\colon \xi\in \Xi(K,G,L)\}$. Let $\K$ be the weak* closure of $\E$.
Then 
\eqn{amc}
$$
AM_c(\Gamma(K,G,L))=AM_c(\Xi(K,G,L))=AM_c(\E)=AM_c(\K).
$$
Only the last equality is not obvious. Let $(\xi_j)$ be a sequence of curves from 
$\Xi(K,G,L)$ such that $\nu_{\xi_j}$ converge weak* to $\nu\in\K$. By the Arzel\`a-Ascoli theorem 
(see \cite[p. 169]{Roy})
there exists a subsequence (not relabelled) which converges uniformly to a limit curve
$\xi$, and, by compactness of 
$K$ and openness of $G$, we have $\xi\in \Xi(K,G,L)$. For each non-negative continuous function $\rho$ on
$X$ we have
$$
\int_{\xi}\rho\,ds\le \liminf_j \int_{\xi_j}\rho\,ds= \lim_j\int_{X}\rho\,d\nu_{\xi_j}=\int_{X}\rho\,d\nu.
$$
It follows that each admissible sequence for $AM_c(\E)$ is also admissible for $AM_c(\K)$ and 
thus $AM_c(\K)\le AM_c(\E)$, whereas the converse inequality is obvious. This proves \eqref{amc}.
By \cite[Theorem 5.5]{HEKMM}, $AM(\K)=M_1(\K)$ (as $\K$ is compact) and 
by \cite[Theorem 3.4]{HEKMM}, $AM=AM_c$.
Hence
$$
\aligned
M_1(\Gamma(K,G,L))&\le M_1(\K)=AM_c(\K)=AM_c(\Gamma(K,G,L))\\&
=AM(\Gamma(K,G,L))\le AM(\Gamma(K,G)).
\endaligned
$$
Passing to the supremum over $L$ we obtain the conclusion.
\end{proof}

\begin{lemma}\label{LSus} If $E \subset G$ is a Suslin set, then 
$\cp_1(E,G) \leq AM(\Gamma(E)).$
\end{lemma}

\begin{proof}
Since $E$ is a Suslin set, Proposition \ref{p:cap}(f) implies that there are compact sets
$K_1 \subset K_2 \subset \, ... \, \subset E$ such that
$\cp_1(E,G) = \lim_i \cp_1(K_i,G)$. Now by Lemma \ref{Lcap} 
$$\cp_1(K_i,G) = AM(\Gamma(K_i,G)) \leq AM(\Gamma(E))$$
because $\Gamma(K_i,G) \subset \Gamma(E)$.
\end{proof}

\begin{lemma} \label{Linc}
Let $K_1 \subset K_2 \subset \, ... \, $ be compact sets in G with
\begin{equation} \label{Ecapfin}
\lim_{i \rightarrow \infty}\cp_1(K_i,G) < \infty.
\end{equation}
Then there is a $BV$ function $w$ in $X$ such that $w = 0$ in $X \setminus G$,
$w =1$ on $\bigcup_i K_i$, $0 \leq w \leq 1$ and 
\begin{equation} \label{Ecapfin1}
V(w,X) \leq \lim_{i \rightarrow \infty} \cp_1(K_i,G).
\end{equation}
\end{lemma}

\begin{proof}
For each $i$ pick $u_i \in \Lip_0(K_i,G)$ such that $0\le u_i\le 1$ 
and 
$$ 
\int_G |\nabla u_i|\, d\mu \leq \cp_1(K_i,G) + 1/i.
$$
By the compact embedding of $BV$ into $L^1_{\loc}$, see  \cite[Theorem 
3.7]{Mi}, there is a limit function $w$ and a subsequence $(v_i)_i$ of $(u_i)_i$ such that
$v_i\to w$ in $L_{\loc}^1(X)$ and $\mu$-a.e. In particular, we can assume that $w=1$ 
on $\bigcup_i K_i$ and \eqref{Ecapfin1} holds.
\end{proof}

We recall some measure theoretic notation. Let $E \subset X$ be a ($\mu$--) measurable set. The \textit{measure
theoretic boundary} $\partial_*E$ of $E$ consists of points
$x \in X$ such that
$ \Theta(x,E) > 0$
and $\Theta(x,X \setminus E) > 0$
where
$$\Theta(x,A) = \limsup_{r \rightarrow 0}\frac{\mu(B(x,r) \cap A)}{\mu(B(x,r))} $$
is the upper $\mu$--density of $A$ at $x$.
 The \textit{measure
theoretic interior} $\inte E$ 
and the \textit{measure theoretic exterior} $\exte E$ of $E$
are the sets
of points $x \in  X$ where $\Theta(x,X \setminus E) =0$ and
$\Theta(x,E)=0$, respectively. 
The sets $\partial_*E$, $\inte E$ and $\exte E$ are Borel sets.

For an open bounded set $G \neq X$ and $E \subset G$ we define the \textit{perimeter capacity} of $E$ in 
$G$ as 
$$\Cp(E,G) = \inf\Big\{P(F,X) \colon E \subset \inte F, \,F \subset G \textnormal{ measurable} \Big\}.$$
Note that the perimeter of $F$ is relative to $X$ and not relative to $G$.

\begin{lemma}\label{cap=Cap}
If $E$ is a Suslin set in $G \subset X$ and $AM(\Gamma(E)) < \infty$, then 
\begin{equation} \label{ECc1}
 \Cp(E,G) \le \,\cp_1(E,G).
\end{equation}
\end{lemma}

\begin{proof}
Let $U$ be an open set such that $E \subset U \subset G$.
By Lemma \ref{LSus} we have $\cp_1(E,G) < \infty$. 
Next choose compact sets $K_1 \subset K_2 \subset \, ... \subset U$ such that 
$\bigcup_i K_i = U$; now 
$$ \cp_1(K_i,G) \leq \cp_1(U, G)$$ 
for all $i$. 

Let $w$ be the $BV$ function in Lemma \ref{Linc}. Note that $w = 1$ in $U = \bigcup_i K_i$.
By the co--area formula \cite[Proposition 4.2]{Mi}
and Lemma \ref{Linc}
$$\int_0^1P(\{x: w(x) > t \}, X) \,dt \le V(w,X) \leq \lim_i \cp_1(K_i, G) \leq \cp_1(U, G).$$
 Thus there
is some $t \in (0,1)$ such that the set $A =\{x: w(x) > t \}$
has finite perimeter, $\inte A \supset E$ and
$P(A, X) \leq \cp_1(U, G)$. Note that it is possible that $A = G$. Since 
$$\Cp(E,G) \leq P(A,X) \leq  \cp_1(U, G)$$
and this holds for all open sets $U$ with $E \subset U \subset G$ we obtain (\ref{ECc1}).
\end{proof}

\section{$AM(\Gamma(E)) \leq C \, co\H^1(E)$}

Throughout this section we assume that $(X,d)$ and $\mu$ satisfy the assumptions (A).

We need the following auxiliary lemma for the main result. Note that the set $E$ below is an arbitrary subset of $X$.

\begin{lemma}\label{L=0} If $E \subset X$ and $AM(\Gamma(E)) < \infty$, then $\mu(E) = 0$.
\end{lemma}

\begin{proof}
By \cite[Theorem 2]{HEMM2} there is a co--Suslin set $E' \supset E$ such that $AM(\Gamma(E')) = AM(\Gamma(E)$. 
Since co--Suslin sets are $\mu$--measurable we may assume that $E$ is $\mu$ measurable and since we can also assume that $E$ is bounded, it suffices to prove the lemma in the case $\mu(E) < \infty$.

Let $\varepsilon > 0$. Since $\mu(\overline{B}(x,r) \setminus B(x,r))=0$ except for a countable set of $r>0$ we find by the Vitali covering theorem disjoint
balls $\overline{B}(x_i, r_i)$ such that
$r_ i < \varepsilon$ and $\bigcup_i B(x_i, r_i)\supset E \setminus E_0$ where $\mu(E_0) = 0$. Now we can replace $E$ by $E \setminus E_0$ which we continue to denote by $E$.

Fix $B_i = B(x_i, r_i)$ and let $K \subset E \cap B_i$ be compact. For $\delta > 0$ pick $u \in N^{1,1}_0(B_i)$ such that $u = 1$
on $K$, $0 \leq u \leq 1$ and
$$\int_{B_i}g_u \, d\mu < \cp_1(K,B_i) + \delta.$$
By the Poincar\'e inequality \cite[Theorem 5.51]{BB} for $N^{1,1}_0(B_i)$--functions there is a 
constant $C$ depending only on $C_P$ and $C_{\mu}$ so that
$$\mu(K) \leq \int_{B_i}u \, d\mu \leq C r_i 
\int_{B_i} g_u \, d\mu <
C r_i  (\cp_1(K,B_i) + \delta)$$
and letting $\delta \rightarrow 0$ we obtain 
from Lemma \ref{Lcap}
$$\mu(K) \leq C r_i  AM(\Gamma(K,B_i)) \leq
 Cr_i AM(\Gamma(E\cap B_i,B_i)). $$
Since this holds for all compact sets $K \subset E\cap B_i $
$$\mu(E\cap B_i) \leq C r_i \, AM(\Gamma(E\cap B_i,B_i)).$$
The path families $\Gamma(E\cap B_i,B_i)$
lie in the disjoint sets $\overline{B}_i$ and
are subfamilies of $\Gamma(E)$. Summing over $i$ we obtain
$$ \mu(E) = \sum_i \mu(E\cap B_i) \leq C \sum_i
r_i \, AM (\Gamma(E\cap B_i,B_i)) \leq 
C \varepsilon \, AM(\Gamma(E)),$$
and $\varepsilon \rightarrow 0$ completes the proof.
\end{proof}

The comparison of the $BV$ capacity with the $(n-1)$--dimensional 
Hausdorff content is due to Fleming \cite{Fle}. It has been 
generalized to the framework of metric measure spaces by  Kinnunen, 
Korte, Shanmugalingam and Tuominen \cite{KKST1}. Here we need a 
version for the $\delta$-Hausdorff content related to the $co\H^1$--measure.

\begin{lemma}
\label{Lnull}
Let $M$ be a bounded open set in $X$.
For $\delta>0$ there exists
$\alpha>0$ such that for each 
open set
$G$ with $\mu(G)<\alpha$ and $E\subset G\subset M$
we have
\begin{equation} \label{ELnull}
co\H^1_{\delta}(E) \le C\, \Cap(E,G),
\end{equation}
where $C$ depends only on $C_P$, $\lambda_P$ and $C_{\mu}$.
\end{lemma}

\begin{proof} 
We write for $C$ a generic constant which depends only on $C_P$, $\lambda_P$ and $C_{\mu}$.

Set $\delta' = \delta/(5 \lambda_P)$ and $\kappa = 4 C_P$.  
Let $G$ be a bounded open set such that $E \subset G \subset M$. We find
$\alpha > 0$ such that for each $x \in G$
\begin{equation} \label{EG1}
\mu(B(x,\delta') \cap G) \leq \frac{1}{\kappa}
\mu(B(x, \delta'))
\end{equation}
provided that
$\mu(G) < \alpha$. Suppose that no such  $\alpha$ exists. Then there are
open sets $G_i$ and $x_i \in G_i$ such that $E \subset G_i \subset M$ and
$$
\frac{1}{i}> \mu(G_i) \geq \mu(B(x_i,\delta') \cap G_i) > \frac{1}{\kappa}\mu(B(x_i, \delta'))
$$
but because each $x_i$ belongs to a fixed bounded set $M$, $\mu(B(x_i, \delta')) > c > 0$ which
 leads to contradiction.
 
Fix $G$ as above. To prove (\ref{ELnull}) we may assume that $\Cap(E,G) < \infty$ and
for $\varepsilon > 0$ we choose a competitor $F \subset G$ for $\Cp(E,G)$ with
$P(F,X) \leq \Cp(E,G) + \varepsilon$. Let $x \in E$, $B(r) = B(x,r)$ and define
$$r_x = \inf \{r > 0: \, \mu(F \cap B(r)) \leq
 \frac{1}{2C_P} \mu(B(r) \}.$$
 Now $0 < r_x < \delta'$ because
$$ \lim_{r \rightarrow 0} \frac{\mu(F \cap B(r))}{\mu(B(r))} = 1$$
and by (\ref{EG1})
$$\mu(F \cap B(\delta')) \leq \mu(G \cap B(\delta') )
\leq \frac{1}{4C_P}\mu(B(\delta')) <
\frac{1}{2C_P}\mu(B(\delta')).$$

Let $r < r_x$. Then
$$\mu(F \cap B(r_x)) \geq \mu(F  \cap B(r))
> \frac{1}{2C_P} \mu(B(r))$$
and letting $r \rightarrow r_x$ we obtain
\begin{equation} \label{EG2}
\mu(F \cap B(r_x)) \geq \frac{1}{2C_P} \mu(B(r_x)).
\end{equation}
On the other hand we show that
\begin{equation} \label{EG3}
\mu(F \cap B(r_x)) \leq \frac{1}{2} \mu(B(r_x)).
\end{equation}
If $\mu(F \cap B(r_x)) \leq  \mu(B(r_x))/(2C_P)$, then equality holds 
in (\ref{EG2}) and (\ref{EG3}) is immediate. If
$$\mu(F \cap B(r_x)) > \frac{1}{2C_P} \mu(B(r_x))$$
then by the definition of $r_x$ there is
$r \in (r_x, 2r_x)$ such that
$$\mu(F \cap B(r_x)) \leq \mu(F \cap B(r)) \leq \frac{1}{2C_P} \mu(B(r)) \leq \frac{1}{2} \mu(B(r_x)).$$

Next we use the $BV$--Poincar\'e inequality (\ref{EPBV}) for the $BV$ function $\chi_F$ in $B(r_x)$. 
By (\ref{EG2}) and (\ref{EG3})
$$\frac{1}{2C_{\mu}} \leq (\chi_F)_{B(r_x)} = \frac{\mu(F \cap B(r_x))}{\mu(B(r_x))} \leq \frac{1}{2} $$
and we obtain
$$
\aligned
\frac{\mu(B(r_x))}{4 C_{\mu}}& \leq 
\frac{ \mu(F \cap B(r_x))}{2} \leq \int_{F \cap B(r_x)} (1-  (\chi_F)_{B(r_x)} ) \, d\mu
\\
&\leq \int_{B(r_x)} |(\chi_F-  (\chi_F)_{B(r_x)} | \, d\mu \leq C_P \, r_x P(  F, B(\lambda_P \, r_x))
\endaligned$$
and so
\begin{equation} \label{EG4}
\frac{\mu(B(r_x))}{r_x} \leq C \, P(  F, B(\lambda_P r_x)).
\end{equation}
By the $5$--covering lemma we find balls $B_j = B(x_j, \lambda_P r_{x_j})$ from 
the collection $\{ B(x,\lambda_P r_x) \}$ so that 
the balls $B_j$ are disjoint and the balls $5B_j = B(x_j, 5\lambda_P r_{x_j})$ cover
$E$. Set $D = \bigcup_j 5 \,B_j$. Since $ 5 \lambda_P r_{x_j} < 5\, \lambda_P \delta' = \delta$ 
we obtain from (\ref{EG4})
$$ co\H^1_{\delta}(E)  \leq
\sum_j \frac{\mu(5B_j)}{5r_{x_j}} \leq
C \sum_j \frac{\mu(B(x_j, r_{x_j}))}{r_{x_j}}$$
$$\leq C \, \sum_j P(F, B_j)
\leq  C \, P(F, X) \leq C (\Cap(E,G) + \varepsilon)$$
where the doubling property of $\mu$  and the fact that the balls $B_j$ are disjoint have also been used.
Letting $\varepsilon \rightarrow 0$ we complete the proof.
\end{proof}

The following lemma combines the achieved results.

\begin{lemma} \label{Lcomp}
Suppose that $E \subset X$ is a bounded Suslin set such that $AM(\Gamma(E)) < \infty$. Then
\begin{equation} \label{Ecompp}
co\H^1(E) \le C \, AM(\Gamma(E)) 
\end{equation}
where the constant $C$ depends only on $C_P, \, \lambda_P$ and $C_{\mu}$.
\end{lemma}

\begin{proof}
Lemma \ref{L=0} yields $\mu(E) = 0$. Fix $\delta > 0$ and then, by Lemma \ref{Lnull}, we 
find a bounded open set $G \neq X$ containing $E$ with
$$co\H^1_{\delta}(E)\le C \, \Cp(E,G).$$
Now Lemmata \ref{cap=Cap} and
\ref{LSus} imply
$$
\Cp(E,G) \leq \cp_1(E,G) \leq AM(\Gamma(E))$$
and hence $co\H^1_{\delta}(E) \le C \, AM(\Gamma(E)).$ Passing to 
the supremum w.r.t. $\delta>0$ we obtain (\ref{Ecompp}).
\end{proof}

\begin{thm} \label{Tsigma}
Let  $E \subset X$ be a Suslin set. Then
\begin{equation} \label{Ecomp}
C_1 \, co\H^1(E) \le  \, AM(\Gamma(E)) \leq C_2 \, co\H^1(E)
\end{equation}
where the constant $C_1 > 0$ depends only on $C_P, \, \lambda_P$ and $C_{\mu}$ 
and the constant $C_2$ only on $C_{\mu}$.
\end{thm}

\begin{proof}
The second inequality in (\ref{Ecomp}) follows from Theorem \ref{Tupper}. For the first inequality 
fix $x_0\in X$ and observe that
$$
C_1 co\H^1(E\cap B(x_0,j))\le AM(\Gamma(E\cap B(x_0,j))) \le  
AM(\Gamma(E)),\qquad j=1,2,\dots
$$
by Lemma \ref{Lcomp}.
Letting $j\to \infty$ we conclude the proof.
\end{proof}

If $E \subset X$ has $\sigma$--finite $co\H^1$--measure,
then Theorem \ref{Tsigma} holds without the assumption that 
$E$ is a Suslin set.

\begin{thm} \label{Tsigma1}
Suppose that $E \subset X$ has $\sigma$--finite $co\H^1$--measure. Then 
\begin{equation} \label{Ecomp1}
C_1 \, co\H^1(E) \le  \, AM(\Gamma(E)) \leq C_2 \, co\H^1(E)
\end{equation}
where the constants $C_1 $ and $C_2$ are as in Theorem \ref{Tsigma}.
\end{thm}

\begin{proof}
The right inequality of (\ref{Ecomp1}) again follows from Theorem
\ref{Tupper}. 
For the left inequality suppose first that $co\H^1(E)<\infty$. Then 
there is a Borel set $F\supset E$ such that $co\H^1(F)=co\H^1(E)$ and a 
co-Suslin set $E'\supset E$ such that $AM(\Gamma(E'))=AM(\Gamma(E))$,
see \cite[Theorem 2]{HEMM2}. We may assume that $E'\subset F$.
Then the set function 
$$
\nu\colon A\mapsto co\H^1(A\cap F),\qquad A\text{ Borel}
$$
is a finite Borel measure. We extend $\nu$ to 
the class of all $\nu$-measurable sets by completion. Then the set
$E'$ is $\nu$-measurable as it is co-Suslin \cite[Theorem 21.10]{Kech}. 
It follows that there
is a Borel set $A\subset E'$ such that $\nu(A)=\nu(E')$
\cite[Theorem 17.10]{Kech}.
Now, 
$$
co\H^1(E)\le co\H^1(E')=\nu(E')=\nu(A)=co\H^1(A)
$$
and 
$$
AM(\Gamma(A))\le AM(\Gamma(E'))=AM(\Gamma(E)).
$$
Since $C_1\,co\H^1(A)\le AM(\Gamma(A))$, we conclude that
$$
C_1\,co\H^1(E)\le AM(\Gamma(E)).
$$

In the general case we find $E_1\subset E_2\subset\dots$ such that 
$co\H^1(E_i)<\infty$ and $E=\bigcup_i E_i$. Let 
$F_i$ be Borel set such that $F_i\supset E_i$ and 
$co\H^1(F_i)=co\H^1(E_i)$.
Since $E_1\subset F_1\cap F_2\subset F_1$, we have 
$co\H^1 (F_1\setminus F_2)=co\H^1(F_1)-co\H^1 (F_1\cap F_2)=0$
and thus $co\H^1(F_1\cup F_2)\le co\H^1(E_2)$. Continuing by induction we 
may assume that $F_1\subset F_2 \subset \dots$. Therefore
$$
co\H^1(E)\le co\H^1\Big(\bigcup_iF_i\Big)= \lim_i co\H^1(F_i)= \lim_i 
co\H^1(E_i)\le C_1^{-1}AM(\Gamma(E)).
$$
\end{proof}

In the Euclidean setting, the $co\H^1$ measure satisfies
$$\alpha_{n-1}co\H^{1}(E)=\alpha_n\haus(E),$$ 
where
$$\haus(E) = \sup_{\delta> 0}\haus_{\delta}(E)$$
is the spherical Hausdorff measure defined through the 
spherical Hausdorff $\delta$-content
$$
 \haus_{\delta}(E)= \inf \Bigl\{
 \sum_{i=1}^{\infty} \alpha_{n-1}r_i^{n-1} \colon\,
 E \subset \bigcup_{i=1}^{\infty} B(x_i,r_i)\,, \, r_i<\delta \Bigr\}
 $$
and
$\alpha_m$ denotes the volume of the $m$-dimensional unit ball.
It is easily seen that the spherical Hausdorff measure is equivalent to the standard Hausdorff
measure $\standhaus$ defined in terms of diameters, namely
$$
\standhaus(E)\le \haus(E)\le 2^n\standhaus(E),\qquad E\subset\rn,
$$
see \cite[2.10.2]{Fed}.
Now, Theorems \ref{Tsigma} and  \ref{Tsigma1} 
yield (with properly modified constants):

\begin{cor} \label{Crn}
If $E$ is a Suslin set in $\rn$ or has $\sigma$--finite $\standhaus$--measure, then
$$C_1 \, \standhaus(E) \leq AM(\Gamma(E)) \leq C_2 \, \standhaus(E)$$
where the positive constants $C_1$ and $C_2$ depend only
on $n$.
\end{cor}

\section{perimeter
 and  $AM$--modulus in $X$}

We  characterize sets $E$ of finite perimeter in $X$ using the $AM$--modulus
of the path family  $\Gamma(\partial_*E)$.
Such a characterization was presented for $X= \rn$  in \cite{HEMM2}. 

We also study the connection of the perimeter of $E$ in an open set $\Omega \subset X$ to the family $\Gamma_{\mathrm{cross}}(E, \Omega)$ whose paths lie in an open set $\Omega$ and  meet both the measure theoretic exterior and interior
of $E$ and present a measure theoretic version of the elementary topological fact. Namely, if $X$ is a topological space, $E \subset X$ and int$\, E$, ext$\,E$ and $\partial E$ are the (topological) interior, exterior and boundary of $E$, respectively, then every curve $\gamma:[a,b] \rightarrow X$ which meets int$\, E$ and ext$ \,E$ also meets $\partial E$. We show that $AM$ a.e. path 
$\gamma \in \Gamma_{\mathrm{cross}}(E, \Omega)$ meets the measure theoretic boundary $\partial_*E$ of $E$ provided that $E$ has finite perimeter in $\Omega$. In \cite[Theorem 5.3]{KL} a closely related result is
proved under more restrictive assumptions on $E$ for the $M_1$--modulus. 

We assume that $X$ satisfies (A) and, as before, $C$ is a constant 
which depends only on $C_{\mu}$, $C_{\lambda}$ and $C_P$ and can change inside a line. 

\begin{lemma} \label{LPAM}
If $\Omega$ be an open set in $X$ and $E \subset X$ measurable, then
$$AM(\Gamma_{\mathrm{cross}}(E, \Omega)) \leq C\,P(E,\Omega).$$
\end{lemma}

\begin{proof}
Let $u$ be the Lebesgue representative of $\chi_E$,
i.e. 
$$u(x) = \lim_{r \rightarrow 0}\frac{\mu(E \cap B(x,r))}{\mu(B(x,r))}$$
whenever the limit exists, then $u(x) = 1, \, x \in \inte E$, $u(x) = 0, \, x \in \exte E$ and $u = \chi_E$ a.e. in $\Omega$.

For the proof we may assume that $P(E, \Omega) < \infty$ and then we can use the special sequence of locally
Lipschitz functions constructed in \cite[Proposition 4.1]{KKST}; i.e. there is a sequence $u_k \in \textnormal{Lip}_{\textnormal{loc}}(\Omega)$ such that
$u_k \rightarrow u$ pointwise $co\H^1$ a.e. in $\Omega \setminus \partial_*E$, $u_k \rightarrow u$
in $L^1(\Omega)$ and
\begin{equation} \label{EPAM1}
\liminf_{k \rightarrow \infty} \int_{\Omega}|\nabla u_k| \, d\mu \leq C \, P(E, \Omega).
\end{equation}

Let $A \subset \Omega \setminus \partial_*E$ be the set where $\lim_k u_k(x) \neq u(x)$. Now
$co\H^1(A) = 0$ and by Theorem \ref{Tupper},
$AM(\Gamma(A)) = 0$.  The sequence of functions $|\nabla u_k|$ is $AM$--admissible for
$\Gamma_{\mathrm{cross}}(E, \Omega) \setminus \Gamma(A)$
since if $\gamma \in
\Gamma_{\mathrm{cross}}(E, \Omega) \setminus \Gamma(A)$ then there are points
$t_1, \, t_2 \in [0, \ell]$ such that $\gamma(t_1)
\in \inte E$, $\gamma(t_2) \in \exte E$ and
$$1 = \lim_{k \rightarrow \infty}|u_k(\gamma(t_1)) -
u_k(\gamma(t_2))| \leq \liminf_{k \rightarrow \infty}\int_{\gamma} |\nabla u_k| \,ds.$$
By (\ref{EPAM1})
$$AM(\Gamma_{\mathrm{cross}}(E, \Omega)\setminus \Gamma(A)) \leq \liminf_{k \rightarrow \infty} \int_{\Omega} |\nabla u_k| \,d\mu \leq C\,P(E,\Omega)$$
and since $AM(\Gamma(A)) = 0$ we have
$$AM(\Gamma_{\mathrm{cross}}(E, \Omega)) \leq C \, P(E,\Omega).$$
\end{proof}

\begin{thm} \label{Tomit}
If $P(E,\Omega) < \infty$ then $AM$ a.e. path $\gamma \in \Gamma_{\mathrm{cross}}(E,\Omega)$ meets $\partial_*E$.
\end{thm}

\begin{proof}
Let $\Gamma$ be the family of paths in $\Gamma_{\mathrm{cross}}(E, \Omega)$ which do not meet
$\partial_*E$.
By the subadditivity of the $AM$--modulus we may assume
that $\Omega$ is bounded. By \cite[Theorem 4.4 and Theorem 4.6]{AMP} for every open
set $G \subset \Omega$
$$ P(E,G) = \int_{\partial_*E \cap G}\theta \, dco\H^1$$
where $\theta = \theta_E$ is a Borel function
with $1/C \leq \theta \leq C$ in $\Omega$
and, moreover, $co\H^1(\partial_*E \cap \Omega)< \infty$.
Let $\varepsilon > 0$. Now we find a compact set $K \subset \partial_*E \cap \Omega$ such that $P(E,G) < \varepsilon$ for $G = \Omega \setminus K$.

Next observe that $\Gamma \subset \Gamma_{\mathrm{cross}}(E, G)$ because each $\gamma \in \Gamma$ does not meet $K$. By Lemma \ref{LPAM}
$$AM(\Gamma ) \leq AM( \Gamma_{\mathrm{cross}}(E, G))  \leq C \, P(E,G) \leq C \, \varepsilon$$
and letting $\varepsilon \rightarrow 0$ we complete the
proof.
\end{proof}

\begin{thm} \label{TAhlPAM}
Suppose that  $E \subset X$ is a ($\mu$--) measurable set. Then for each open set $\Omega \subset X$ 
\begin{equation} \label{EAPAM}
C_1 P(E, \Omega) \leq AM(\Gamma(\partial_*E \cap \Omega)) \leq C_2 P(E, \Omega)
\end{equation}
where the constants $C_1$ and $C_2$ depend only
on $C_P$, $C_{\lambda}$ and $C_{\mu}$.
\end{thm}

\begin{proof}
For the right inequality in (\ref{EAPAM}) we may assume that 
$P(E, \Omega) < \infty$ and
then by \cite[Theorem 4.4]{AMP}, $$co\H^1(\partial_*E \cap \Omega) \leq C \, P(E, \Omega)$$
and now Theorem \ref{Tupper} gives the required inequality. 

For the left side of (\ref{EAPAM}) we note that  $\partial_*E \cap \Omega$ is a Borel set and thus Theorem  \ref{Tsigma} yields
$$ co\H^1(\partial_*E \cap \Omega)) \leq
C \, AM(\Gamma(\partial_*E \cap \Omega)) < \infty .$$
By the recent result of Lahti \cite[Theorem 1.1]{La} this implies that
$P(E, \Omega) < \infty$ and we can  apply again  \cite[Theorem 4.4]{AMP} to conclude 
$$P(E, \Omega) \leq C \, co\H^1(\partial_*E \cap \Omega)) $$
and complete the proof.
\end{proof}

\section{Geometry of level sets in $X$}

The results in the previous sections can be used to study the structure of level sets of $BV$ 
and continuous functions in $X$ and the latter case together with the results in Section 4
produces a plenitude of open sets in $X$ with $co\H^1$ finite boundaries.

We assume that $X$ satisfies the hypotheses (A) and recall some measure theoretic concepts
asociated with $BV$--functions.

For a measurable set $E$ and $x \in X$ we let
$$\overline{D}(E,x) = \limsup_{r \rightarrow 0}\frac{\mu(E \cap B(x,r))}{\mu(B(x,r))},\,
\underline{D}(E,x) = \liminf_{r \rightarrow 0}\frac{\mu(E \cap B(x,r))}{\mu(B(x,r))},$$
and $D(E,x) = \overline{D}(E,x)$ if $\overline{D}(E,x) = \underline{D}(E,x)$.

Let $\Omega$ be an open set in $X$ and $u\in BV(\Omega)$. The upper 
and lower \textit{approximate limits} of $u$ at $x \in \Omega$
are
$$ u^+(x) = \inf\{s: \, D(\{u >s \},x) = 0\} \textnormal{ and }
u^-(x) = \sup\{t: \, D(\{u <t \},x) = 0\}.$$
Then it is immediate that $u^-(x)\le u^+(x)$.
The function $u$ is \textit{approximately continuous} at $x$ if
$u^+(x) = u^-(x)=u(x)$. This holds
a.e.\ in $\Omega$ by the Lebesgue differentiation theorem. 
The set $J_u = \{u^- < u^+ \}$ is called the jump set of $u$
and it has zero $\mu$--measure, see \cite{KKST}.

For $-\infty\le s,t,\le \infty$ we consider the measure theoretic level sets of
$u \in BV(\Omega)$
$$
\aligned
E^t&=\{x\in\Omega\colon u^-(x)\le t\},\\
E_s&=\{x\in\Omega\colon u^+(x)\ge s\},\\
E_s^t&=E_s\cap E^t,\\
\Lambda_t &= E_t^t.
\endaligned
$$

\begin{lemma} \label{Lorder}
If $u \in BV(\Omega)$, then
\begin{equation} \label{Eorder1}
\mu( \Lambda_t) = 0, 
\end{equation}
and consequently $P(E_t,\Omega)=P(E^t,\Omega)$,
for a.e. $t \in \er$.

 If $u$ is (approximately) continuous at $x$, then
$x\in \Lambda_{u(x)}$.
\end{lemma}

\begin{proof}
To prove (\ref{Eorder1}) note that 
$\Lambda_t \subset A_t \cup J_u$,
where
$$A_t = \{ x\in\Omega\colon \, t = u^{-}(x) = u^{+}(x)\}.$$
Since $A_t \cap A_{t'} = \emptyset$ for $t \neq t'$  and $\mu(J_u) =0$, (\ref{Eorder1})
follows. If $\mu( \Lambda_t) = 0$, then $E_t$ differs from $\Omega\setminus E^t$
by a $\mu$-null set and thus $P(E_t,\Omega)=P(E^t,\Omega)$.

If $u$ is approximately continuous at $x$ and $t=u(x)$, then
$t=u^+(x)=u^-(x)$ and thus $x\in \Lambda_t$.
\end{proof}

\begin{thm} \label{Tlevelper}
Let $u\in BV(\Omega)$. Then for a.e.\ $t\in \er$ we have
\begin{equation} \label{Elevel}
co\H^1(\Lambda_t)\le C\, P(E^t, \Omega)
\end{equation}
where $C$ depends only
on $C_P$, $C_{\lambda}$ and $C_{\mu}$.
\end{thm}

\begin{proof}
We first assume that $\Omega$ is bounded.
Let $T$ be the essential infimum of $u$. Then \eqref{Elevel}
obviously holds for $t<T$. If $t>T$, then $\mu(E^t)>0$
and then also $P(E^t, \Omega)>0$ by the isoperimetric inequality
(see e.g.\ \cite{KL}).
Denote $\psi(t)=P(E^t, \Omega)$
and note that $\psi$ is integrable, see \cite{A2} and
\cite{Mi}.
Let $\tau>T$ be a Lebesgue point for $\psi$ such that 
$\mu(\Lambda_\tau)=0$. By Lemma \ref{Lorder} and the Lebesgue differentiation theorem,
a.e.\ $\tau>T$ has these properties.
We show that $t=\tau$
has the required property. 
Choose $\delta>0$.  Lemma \ref{Lnull} gives $\alpha > 0$ such that for each bounded
open set  
$G$ with $\mu(G)< \alpha$
and $E \subset G$ we have
$$co\H^1_{\delta}(E) \leq C\, \Cp(E,G).$$

Now, using Lemma \ref{Lorder} we find
$a,b\in\er$ such that $a<\tau<b$,
$\psi(a)\le 2\psi(\tau)$, $\psi(b)\le 2\psi(\tau)$,
$\mu(\Lambda_a)=\mu(\Lambda_b)=0$ 
and $\mu(E_a^b)<\alpha$. We find an open
set $G \supset E_a^b$ such that still
$\mu(G) < \alpha$. 
Choose $x\in \Lambda_{\tau}$. Then $a<u^+(x)$, $u^-(x)<b$, and thus
$x\in \partial_*E_a$ (if $\overline D(E^a,x)>0$), or 
$x\in \partial_*E^b$ (if $\overline D(E_b,x)>0$), or 
$x\in \inte E_a^b$ (if $D(E^a,x)=D(E_b,x)=0$).
Summarizing,
$$
\Lambda_{\tau}\subset \partial_*E_a\cup \partial_*E^b\cup \inte E_a^b.
$$
We have
$$
\aligned
co\H^1_{\delta}(\partial_*E_a)&\le C\,P(E_a,\Omega)=C\,P(E^a,\Omega)\le 2C \, P(E^{\tau},\Omega),
\\
co\H^1_{\delta}(\partial_*E^b)&\le C \, P(E^b,\Omega))\le 2CP(E^{\tau},\Omega)
\endaligned
$$
and then
$$
\aligned
co\H^1_{\delta}(\inte E_a^b)&\le C\, \Cap(\inte E_a^b,G) \le  C\,P(E_a^b,G)
\\&
\le C(P(E^a,\Omega)+P(E^b, \Omega))\le 4C \, P(E^{\tau}, \Omega).
\endaligned
$$
Letting $\delta \rightarrow 0$ we obtain (\ref{Elevel}).

Suppose that $\Omega$ is unbounded. Fix a point $x_0 \in X$ and for each $i = 1,\,2, \, ...$ let 
$\Omega_i = \Omega \cap B(x_0, i)$ and $u_i = u|\Omega_i$. Denote by $E^t(u_i)$ the set $E^t$ associated
with $u_i$ and other sets, like $\Lambda_{\tau}(u_i)$, similarly. Now for a.e. $t \in \er$,
$\mu(\Lambda_t(u_i))=0$ for every $i$ and so for a.e. $t \in \er$
$$co\H^1(\Lambda_{t}(u_i)) \leq C\, P(E^t(u_i), \Omega_i) \leq C \,P(E^t(u_i), \Omega)
\le C\, P(E^{t}, \Omega) $$
for every $i$
and this easily implies (\ref{Elevel}) for $u$.
\end{proof}

If $u \in BV(\Omega)$ then by the co--area formula \cite[Proposition 4.2]{Mi} for the perimeter 
$P(E^t, \Omega) < \infty$ for
a.e. $t \in \er$. Hence Theorem \ref{Tlevelper} and Lemma \ref{Lorder} yield

\begin{cor} \label{Clevel}
If $u \in BV(\Omega)$, then 
$$
co\H^1(\Lambda_t)<\infty\text{ for a.e. }t\in\er.
$$
If, in addition,
$u$ is (approximately) continuous, then $co\H^1(u^{-1}(t)) < \infty$ for a.e. $t \in \er$.
\end{cor}

\begin{examples} \label{Exlevel}
The above corollary can be used to construct sets in $X$ whose boundaries have finite $co\H^1$--measure.
For example, let  $u \in BV(\Omega) \cap C(\Omega)$. Then for a.e. $t \in \er$ the boundary
of the open set $\{u > t\}$ has finite $co\H^1$--measure. For a more
specific example let $x_0 \in X$ and take $u(x) = d(x, x_0)$.
It follows that 
the topological boundary $\partial B(x_0,r)$ of the ball $B(x_0,r)$ has 
finite $co\H^1$--measure for a.e. $r > 0$. 
This is an improvement of the earlier results since it has been only known that $\mu(\partial B(x_0,r)) = 0$
except for a countable number of $r$ and that $co\H^1(\partial_* B(x_0,r)) < \infty$ for
a.e. $r>0$. More generally, if $K \subset X$ is a bounded set, then $u(x) = \operatorname{dist}(x, K)$ is a Lipschitz function and thus the boundary of 
the $t$--inflation $\{x :\, \dist(x,K) < t\}$  of $K$ has
finite $co\H^1$--measure for a.e. $t>0$.
\end{examples}


\end{document}